The sum of seven consecutive integers is 980. How many of them are prime? Name the prime numbers.
Found 2 solutions by checkley71, Adam:
Answer by checkley71(8403)

x+(x+1)+(x+2)+(x+3)+(x+4)+(x+5)+(x+6)=980
7x+21=980
7x=980-21
7x=959
x=959/7
x=137 the smallest of these sequential numbers. also a prime number.
137+2=139 also a prime number.
no other prime numbers between 137 & 143.

Answer by Adam(64)

seven consecutive integers - if we take 980/7 = 140, we can suppose that 140 is middle value and probably even middle term, that means that there are 3 on left and 3 on right- whole sequence is thus 137,138,139,140,141,142,143 every of theese numbers have to be checked for beiing primal 137 - yes, 138 can be divided by 2, 139 - yes, 140- i.e. 2, 141 - 3, 142 - 2,143 - 11
two of them are primal = 137 and 139